OpenStudy (anonymous):

Which shows two and two-eighths as a simplified improper fraction? Answer sixteen-eighths nine-fourths eighteen-eighths five-fourths

OpenStudy (lgbasallote):

hint: \[\huge 2 \frac 28 \implies \frac{(8 \times 2) + 2 }{8}\] does that help?